# Markov spin models for image generation : explicit large deviations with respect to the number of pixels

### v1

@@ Abstract
For the discrete-time or the continuous-time Markov spin models for image generation when each pixel $n=1,..,N$ can take only two values $S_{n}=\pm 1$, the finite-time forward propagator depends on the initial and final configurations only via their overlap. The large deviations properties with respect to the number $N$ of pixels are analyzed for the overlap and/or the magnetization during the forward noising dynamics and during the backward reconstructive dynamics for various initial conditions, namely (i) a single image (ii) a mixture of two images (iii) when the initial condition corresponds to the Curie-Weiss mean-field ferromagnetic model.
@@ I Introduction
Besides their algorithmic and practical interests, the generative Markov models (see the review [1] and references therein) are also fascinating in relation with other areas of statistical physics and of stochastic processes, as stressed from various perspectives in the recent works [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 13, 12, 14].
@@
The goal of the present work is to analyze their large deviations properties when the number $N$ of pixels becomes large $N\to+\infty$. Indeed, the theory of large deviations (see the reviews [15, 16, 17] and references therein) has become nowadays the unifying language of both equilibrium and non-equilibrium statistical physics, with remarkable achievements in the field of non-equilibrium stochastic dynamics (see the reviews with various perspectives [18, 19, 20, 21, 22, 23, 24, 25, 26], the PhD Theses [27, 28, 29, 30, 31, 32] and the Habilitation Thesis [33]). In particular, many results have been obtained for time-averaged observables over a long Markov trajectory [34, 18, 23, 24, 25, 26, 28, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 33, 53, 54, 55, 56, 57, 60, 58, 59, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78], or for time-dependent properties in systems involving a large number $N$ of degrees of freedom, either via the Macroscopic Fluctuation Theory (see the review [22] and references therein) or via the formulation in terms of Poisson processes [79, 80, 81, 61, 74].
The field of generative Markov models is dominated by the formulation in terms of diffusion processes, but the case of Markov chains on discrete-state spaces has also been considered [13, 14, 82, 83, 84, 85, 86]. For physicists, it is very natural to focus on spin models where each pixel $n$ can take only two values $S_{n}=\pm 1$ : the simplest continuous-time single-spin-flip model has been recently discussed in [13, 14], but one can also consider its discrete-time counterpart (see section II) that leads to the same form of the forward propagator (see the more detailed discussion in subsection II.3) and thus to identical conclusions for properties concerning finite-time-intervals.
@@
The paper is organized as follows. In section II, we describe the general properties of discrete-time Markov spin models for image generation and the specific form of the forward propagator in terms of the overlap between the initial and the final configurations. In section III, we study the large deviations properties of the forward propagator when the number $N$ of pixels is large $N\to+\infty$. The consequences for the backward-generative dynamics are then analyzed for three cases of initial conditions, namely for a single initial image in section IV, for a mixture of two initial images in section V, and finally for the Curie-Weiss model in section VI. Our conclusions are summarized in section VII.
@@ II Discrete-time Markov spin models for image generation
In this section, we describe the general properties of discrete-time Markov spin models for image generation with $N$ pixels, when each pixel $n=1,..,N$ has only two possible states $S_{n}=\pm 1$, so that there are ${\cal N}=2^{N}$ possible images $\vec{S}=(S_{1},..,S_{N})$. The discrete-time framework is technically clearer to understand the subtleties of the backward-generative dynamics, while the properties concerning finite-time-intervals are the same as in the continuous-time version of the spin model considered in the previous works [13, 14] (see the more detailed discussion in subsection II.3).
@@ II.1 General properties of the forward noising dynamics and of the backward reconstructive dynamics
The probability ${\cal P}^{Traj}[\vec{S}(0\leq t\leq T)]$ to see the forward trajectory $\vec{S}(0\leq t\leq T)$ over the time window $t=0,1,,,,T-1,T$ when the initial condition $\vec{S}(t=0)$ is drawn with the probability $P^{ini}(\vec{S}(0))$
@@
involves the matrix elements $\langle\vec{S}(t)|{\cal W}|\vec{S}(t-1)\rangle$ of the Markov matrix ${\cal W}$ between the image $\vec{S}(t-1)$ at time $(t-1)$ and the image $\vec{S}(t)$ at time $t$. The specific choice of ${\cal W}$ will be discussed later in subsection II.2, while here we describe the general principles.
@@
The summation over all the spins except the configuration $\vec{S}(t)$ at time $t$ produces the probability distribution $P_{t}(\vec{S}(t))$ of the configuration $\vec{S}(t)$ at time $t$
@@
$=\sum_{\vec{S}(0)}\langle\vec{S}(t)|{\cal W}^{t}|\vec{S}(0)\rangle=\langle\vec{S}(t)|{\cal W}^{t}|P^{ini}\rangle P^{ini}(\vec{S}(0))$
@@
where the matrix elements $\langle\vec{S}(t)|{\cal W}^{B}_{t+1/2}|\vec{S}(t+1)\rangle$ of the time-dependent backward Markov generator ${\cal W}^{B}_{t+1/2}$ read in terms of the forward generator ${\cal W}$ and of the initial condition $P^{ini}$
$\langle\vec{S}(t)|{\cal W}^{B}_{t+1/2}|\vec{S}(t+1)\rangle$ $\equiv\frac{\langle\vec{S}(t+1)|{\cal W}|\vec{S}(t)\rangle P_{t}(\vec{S}(t))}{P_{t+1}(\vec{S}(t+1))}=\frac{\langle\vec{S}(t+1)|{\cal W}|\vec{S}(t)\rangle\langle\vec{S}(t)|{\cal W}^{t}|P^{ini}\rangle}{\langle\vec{S}(t+1)|{\cal W}^{t+1}|P^{ini}\rangle}$ (6)
Since this time-dependent backward Markov generator ${\cal W}^{B}_{t+1/2}$ has been defined in order to reproduce the full trajectories probabilities of Eq. 5, it will also automatically reproduce all the marginal properties that can be obtained via summations over some variables in Eq. 5, with the following important consequence for the backward propagator over any finite-time interval.
@@
$=\sum_{\vec{S}(t_{1}<\tau<t_{2})}\prod_{t=t_{1}}^{t_{2}-1}\left[\frac{\langle\vec{S}(t+1)|{\cal W}|\vec{S}(t)\rangle P_{t}(\vec{S}(t))}{P_{t+1}(\vec{S}(t+1))}\right]$
$=\frac{\langle\vec{S}(t_{2})|{\cal W}^{(t_{2}-t_{1})}|\vec{S}(t_{1})\rangle P_{t_{1}}(\vec{S}(t_{1}))}{P_{t_{2}}(\vec{S}(t_{2}))}=\frac{\langle\vec{S}(t_{2})|{\cal W}^{(t_{2}-t_{1})}|\vec{S}(t_{1})\rangle\langle\vec{S}(t_{1})|{\cal W}^{t_{1}}|P^{ini}\rangle}{\langle\vec{S}(t_{2})|{\cal W}^{t_{2}}|P^{ini}\rangle}$
can be considered as the direct generalisation of Eq. 6 concerning a single time step, while the backward propagator ${\cal B}_{0,T}[\vec{S}(0)|\vec{S}(T)]$ for the special case $t_{1}=0$ and $t_{2}=T$
${\cal B}_{0,T}[\vec{S}(0)|\vec{S}(T)]=\frac{\langle\vec{S}(T)|{\cal W}^{T}|\vec{S}(0)\rangle P^{ini}(\vec{S}(0))}{\langle\vec{S}(T)|{\cal W}^{T}|P^{ini}\rangle}=\frac{\langle\vec{S}(T)|{\cal W}^{T}|\vec{S}(0)\rangle P^{ini}(\vec{S}(0))}{\displaystyle\sum_{\vec{\sigma}}\langle\vec{S}(T)|{\cal W}^{T}|\vec{\sigma}\rangle P^{ini}(\vec{\sigma}))}$ (8)
only involves the forward propagator $\langle\vec{S}(T)|{\cal W}^{T}|\vec{\sigma}\rangle$ on the time-window $T$ and the initial condition $P^{ini}(.)$.
@@
The main idea of generative Markov models is to consider the trajectories that are generated by the backward dynamics when in the trajectories probabilities of Eq. 5, the true forward solution $P_{T}(\vec{S}(T))$ at time $T$ is replaced by the steady state $P_{*}(\vec{S}(T))$
@@
The hope is that the slight difference between $P_{*}(\vec{S}(T))$ and $P_{T}(\vec{S}(T))$ for large $T$ will produce probabilities ${\cal P}^{GeneTraj}[\vec{S}(0\leq t\leq T)]$ of generated-backward trajectories that are close to the probabilities ${\cal P}^{Traj}[\vec{S}(0\leq t\leq T)]$ of the forward trajectories, and will lead in particular to images $\vec{S}(0)$ at time $t=0$ with a probability distribution sufficiently close to the initial distribution $P^{ini}(\vec{S}(0))$.
The generated distribution $B^{Gene}_{t}(\vec{S}(t))$ at time $t$ can be obtained from the application of the backward propagator ${\cal B}_{t,T}[\vec{S}(t)|\vec{S}(T)]$ of Eq. 7 to $P_{*}(\vec{S}(T))$
$B^{Gene}_{t}(\vec{S}(t))$ $\equiv\sum_{\vec{S}(T)}{\cal B}_{t,T}[\vec{S}(t)|\vec{S}(T)]P_{*}(\vec{S}(T))=\sum_{\vec{S}(T)}P_{*}(\vec{S}(T))\frac{\langle\vec{S}(T)|{\cal W}^{(T-t)}|\vec{S}(t)\rangle P_{t}(\vec{S}(t))}{P_{T}(\vec{S}(T))}$ (10)
$=\sum_{\vec{S}(T)}P_{*}(\vec{S}(T))\frac{\langle\vec{S}(T)|{\cal W}^{(T-t)}|\vec{S}(t)\rangle\langle\vec{S}(t)|{\cal W}^{t}|P^{ini}\rangle}{\langle\vec{S}(T)|{\cal W}^{T}|P^{ini}\rangle}$
up to the time $t=0$
$B^{Gene}_{0}(\vec{S}(0))$ $=\sum_{\vec{S}(T)}{\cal B}_{0,T}[\vec{S}(0)|\vec{S}(T)]P_{*}(\vec{S}(T))=\sum_{\vec{S}(T)}P_{*}(\vec{S}(T))\frac{\langle\vec{S}(T)|{\cal W}^{T}|\vec{S}(0)\rangle P^{ini}(\vec{S}(0))}{P_{T}(\vec{S}(T))}$ (11)
@@
The comparison between the generated distribution $B^{Gene}_{0}(\vec{S}(0))$ and the initial distribution $P^{ini}(\vec{S}(0))$ as a function of the time-window $T$ can be analyzed via the spectral decomposition of the forward propagator [14].
Now that we have summarized the general principles, let us focus on the explicit choice of the forward dynamics.

### v2

@@ Abstract
For the discrete-time or the continuous-time Markov spin models for image generation when each pixel $n=1,..,N$ can take only two values $S_{n}=\pm 1$, the finite-time forward propagator depends on the initial and on the final configurations of the $N$ spins only via a single global variable, namely the extensive overlap that counts the number of spins that have the same value or not in the two configurations. The joint probability distribution of the overlap and of the magnetization during the forward noising dynamics can be written for any finite number $N$ of pixels and in the limit $N\to+\infty$ to extract the large deviations properties. The consequences for the backward reconstructive dynamics are then analyzed for various initial conditions, namely (i) a single image (ii) a mixture of two images (iii) when the initial condition corresponds to the Curie-Weiss mean-field ferromagnetic model in the microcanonical ensemble, as a simple analog of the manifold-hypothesis concerning continuous generative diffusion models.
@@ I Introduction
Besides their algorithmic and practical interests, the generative Markov models (see the review [1] and references therein) are also fascinating in relation with other areas of statistical physics and of stochastic processes, as stressed from various perspectives in the recent works [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. The forward dynamics transforming incrementally non-trivial images into pure noise is an example of the transient non-equilibrium relaxation towards equilibrium when starting from arbitrary initial conditions, that can be analyzed via the spectral decomposition of the forward propagator [16]. The backward reconstructive dynamics is based on the mathematical possibility of reversing the time-arrow in Markov processes that can be implemented on computers, even it would never naturally occur in real life. Its ability to ’create order out of noise’ might be surprising at first, but can be understood if one thinks of the artificial time-dependent backward generator as some kind of ’stochastic Maxwell demon’ that is sufficiently well informed to drive the system from high entropy towards low entropy. The analysis of this backward reconstructive dynamics actually raises a lot of interesting issues for statistical physicists, in particular in relation with spontaneous symmetry breaking, memorization and glass transitions (see [3, 4, 5, 6, 7, 8, 12, 13] and references therein). So while the practical applications of generative Markov models have recently exploded with very impressive achievements, it seems nevertheless useful to develop a better theoretical understanding from the point of view of statistical physics when the number $N$ of pixels becomes large $N\to+\infty$.
@@
In statistical physics, whenever there is a big parameter, the theory of large deviations (see the reviews [17, 18, 19] and references therein) is used either implicitly (just like Mr Jourdain speaking in prose without knowing it) or explicitly. The explicit use of the language of large deviations has recently played a major role in the unification between the equilibrium part and the non-equilibrium part of statistical physics, and has led to a much deeper understanding of non-equilibrium stochastic processes (see the reviews with various perspectives [20, 21, 22, 23, 24, 25, 26, 27, 28], the PhD Theses [29, 30, 31, 32, 33, 34] and the Habilitation Thesis [35]). In particular, many results have been obtained in two important asymptotic regimes, namely on one hand for time-averaged observables over a single stochastic trajectory observed during a long time-window $T\to+\infty$ [36, 20, 25, 26, 27, 28, 30, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 35, 55, 56, 57, 58, 59, 62, 60, 61, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80], and on the other hand for time-dependent properties in systems involving a large number $N\to+\infty$ of degrees of freedom, either via the Macroscopic Fluctuation Theory for driven diffusive systems (see the review [24] and references therein) or via the formulation in terms of Poisson processes for discrete configuration spaces [81, 82, 83, 63, 76]. In the present work concerning generative Markov models, our goal will be similarly to analyze the time-dependent properties of the forward noising dynamics and of the backward denoising dynamics when the number $N$ of pixels is large $N\to+\infty$.
The field of generative Markov models is dominated by the formulation in terms of diffusion processes, but the case of Markov chains on discrete-state spaces has also been considered [15, 16, 84, 85, 86, 87, 88]. Since computers only work with discrete data, the diffusion models actually need in practice to be discretized both in configuration-space and in time. As a consequence, it seems useful to develop also the field of discrete-state discrete-time generative Markov models in order to have theoretical descriptions in direct correspondence with the computer implementations. The simplest discrete-state framework is when each pixel $n$ can take only two values $S_{n}=\pm 1$ corresponding to the familiar spins of statistical physics and to the black-and-white images of old times; the case of colored images where each pixel can be in $C>2$ different colors would be of course more realistic for practical applications and is also standard in statistical physics from the Potts and clock models, but will not be discussed in the present paper in order to simplify the notations and the discussions. For these generative Markov models involving spins $S_{n}=\pm 1$, the simplest continuous-time single-spin-flip model has been recently analyzed in [15, 16], but one can alternatively consider its discrete-time counterpart as will be discussed in detail in section II. Since these two dynamics actually share exactly the same finite-time forward propagator (see the more detailed discussion in subsection II.3), one obtains the same conclusions for all properties concerning finite-time-intervals. The present work is based on the following key observation: this finite-time forward propagator depends on the initial configuration of the $N$ spins and on the final configuration of the $N$ spins only via a single global variable with a simple physical meaning, namely the extensive overlap $Q$ that counts the number of spins that have the same value or not in the two configurations. It is thus possible to make simple explicit calculations either for a finite number $N$ of pixels or in the limit $N\to+\infty$ to extract the large deviations properties.
@@
The paper is organized as follows. In section II, we describe the general properties of discrete-time generative Markov spin models, as well as the specific form of the forward propagator in terms of the overlap between the initial and the final configurations. In section III, the joint probability distribution of the extensive overlap $Q$ and of the extensive magnetization $M$ during the forward noising dynamics is written for any finite number $N$ of pixels and in the limit $N\to+\infty$ to extract the large deviations properties for the intensive overlap $q=\frac{Q}{N}$ and for the intensive magnetization $m=\frac{M}{N}$. The consequences for the backward-generative dynamics are then analyzed for three cases of initial conditions, namely for a single initial image in section IV, for a mixture of two initial images in section V, and finally for the Curie-Weiss mean-field ferromagnetic model in the microcanonical ensemble in section VI. Our conclusions are summarized in section VII, while some technical details of the large deviations calculations are relegated to Appendices.
@@ II Discrete-time Markov spin models for image generation
In this section, we introduce the notations that will be useful in the whole paper to analyze discrete-time Markov spin models for image generation with $N$ pixels, when each pixel $n=1,..,N$ has only two possible states $S_{n}=\pm 1$, so that there are ${\cal N}=2^{N}$ possible images $\vec{S}=(S_{1},..,S_{N})$.
@@ II.1 General properties of the forward noising dynamics and of the backward reconstructive dynamics
The goal of this subsection is to give a short self-contained summary of the main ideas of Markov generative models within the discrete-time formulation which is more pedagogical to understand the subtleties of the backward-generative dynamics.
@@
When the initial condition $\vec{S}(t=0)$ is drawn with the probability $P^{ini}(\vec{S}(0))$, the probability ${\cal P}^{Traj}[\vec{S}(0\leq t\leq T)]$ of the whole forward trajectory $\vec{S}(0\leq t\leq T)$ during the time window $t=0,1,,,,T-1,T$ simply involves the product over the time $t$ of the matrix elements $\langle\vec{S}(t)|{\cal W}|\vec{S}(t-1)\rangle$ of the Markov matrix ${\cal W}$ between the image $\vec{S}(t-1)$ at time $(t-1)$ and the image $\vec{S}(t)$ at time $t$
@@
The specific choice of ${\cal W}$ will be discussed later in subsection II.2, while here we describe the general principles.
@@
This trajectory probability is the analog of the Feynman path-integral integrand for continuous diffusion models.
The summation of the trajectory probability of Eq. 2 over all the spins except the configuration $\vec{S}(t)$ at time $t$ produces the probability $P_{t}(\vec{S}(t))$ to see the configuration $\vec{S}(t)$ at time $t$ when the initial condition is $P^{ini}$
@@
$=\sum_{\vec{S}(0)}\langle\vec{S}(t)|{\cal W}^{t}|\vec{S}(0)\rangle P^{ini}(\vec{S}(0))=\langle\vec{S}(t)|{\cal W}^{t}|P^{ini}\rangle$
@@
where the matrix elements $\langle\vec{S}(t)|{\cal W}^{B}_{t+1/2}|\vec{S}(t+1)\rangle$ of the time-dependent backward Markov generator ${\cal W}^{B}_{t+1/2}$
$\langle\vec{S}(t)|{\cal W}^{B}_{t+1/2}|\vec{S}(t+1)\rangle$ $\equiv\langle\vec{S}(t+1)|{\cal W}|\vec{S}(t)\rangle\frac{P_{t}(\vec{S}(t))}{P_{t+1}(\vec{S}(t+1))}=\langle\vec{S}(t+1)|{\cal W}|\vec{S}(t)\rangle\frac{\langle\vec{S}(t)|{\cal W}^{t}|P^{ini}\rangle}{\langle\vec{S}(t+1)|{\cal W}^{t+1}|P^{ini}\rangle}$ (6)
involves the reweighting of the forward matrix element $\langle\vec{S}(t+1)|{\cal W}|\vec{S}(t)\rangle$ by the ratio $\frac{\langle\vec{S}(t)|{\cal W}^{t}|P^{ini}\rangle}{\langle\vec{S}(t+1)|{\cal W}^{t+1}|P^{ini}\rangle}$ of two forward propagators starting at the initial condition $P^{ini}$. The logarithm of this ratio is the discrete-space discrete-time analog of the score function $\left(-\vec{\nabla}\ln p_{t}(\vec{x})\right)=\left(-\vec{\nabla}\ln\langle\vec{x}|{\cal W}^{t}|p^{ini}\rangle\right)$ that appears as an effective force in the backward Langevin dynamics of generative diffusion models.
Since this time-dependent backward Markov generator ${\cal W}^{B}_{t+1/2}$ has been defined in order to reproduce the full trajectories probabilities of Eq. 5, it will automatically reproduce all the marginal properties that can be obtained via summations over some variables in Eq. 5, with the following important consequence for the backward propagator over any finite-time interval.
@@
$=\sum_{\vec{S}(t_{1}<\tau<t_{2})}\prod_{t=t_{1}}^{t_{2}-1}\left[\langle\vec{S}(t+1)|{\cal W}|\vec{S}(t)\rangle\frac{P_{t}(\vec{S}(t))}{P_{t+1}(\vec{S}(t+1))}\right]$
$=\langle\vec{S}(t_{2})|{\cal W}^{(t_{2}-t_{1})}|\vec{S}(t_{1})\rangle\frac{P_{t_{1}}(\vec{S}(t_{1}))}{P_{t_{2}}(\vec{S}(t_{2}))}=\frac{\langle\vec{S}(t_{2})|{\cal W}^{(t_{2}-t_{1})}|\vec{S}(t_{1})\rangle\langle\vec{S}(t_{1})|{\cal W}^{t_{1}}|P^{ini}\rangle}{\langle\vec{S}(t_{2})|{\cal W}^{t_{2}}|P^{ini}\rangle}$
involves the three forward propagators $\langle\vec{S}(t_{2})|{\cal W}^{(t_{2}-t_{1})}|\vec{S}(t_{1})\rangle$, $\langle\vec{S}(t_{1})|{\cal W}^{t_{1}}|P^{ini}\rangle$ and $\langle\vec{S}(t_{2})|{\cal W}^{t_{2}}|P^{ini}\rangle$. The expression of Eq. 7 can be considered as the direct generalization of the reweighting formula of Eq. 6 concerning a single time step with $t_{1}=t$ and $t_{2}=t+1$, while the backward propagator ${\cal B}_{0,T}[\vec{S}(0)|\vec{S}(T)]$ for the special case $t_{1}=0$ and $t_{2}=T$ reduces to
${\cal B}_{0,T}[\vec{S}(0)|\vec{S}(T)]=\frac{\langle\vec{S}(T)|{\cal W}^{T}|\vec{S}(0)\rangle P^{ini}(\vec{S}(0))}{\langle\vec{S}(T)|{\cal W}^{T}|P^{ini}\rangle}=\frac{\langle\vec{S}(T)|{\cal W}^{T}|\vec{S}(0)\rangle P^{ini}(\vec{S}(0))}{\displaystyle\sum_{\vec{\sigma}}\langle\vec{S}(T)|{\cal W}^{T}|\vec{\sigma}\rangle P^{ini}(\vec{\sigma})}$ (8)
that only involves the forward propagator $\langle\vec{S}(T)|{\cal W}^{T}|\vec{\sigma}\rangle$ over the time-window $T$ and the initial condition $P^{ini}(.)$.
@@
The main idea of generative Markov models is to consider the trajectories that are generated by the time-dependent backward generator when, in the trajectories probabilities of Eq. 5, the true forward solution $P_{T}(\vec{S}(T))$ at time $T$ is replaced by the steady state $P_{*}(\vec{S}(T))$ of the forward dynamics
@@
So the probability ${\cal P}^{Traj}[\vec{S}(0\leq t\leq T)]$ of the forward trajectory $\vec{S}(0\leq t\leq T)$ is reweighted by the ratio $\frac{P_{*}(\vec{S}(T))}{P_{T}(\vec{S}(T))}$ to obtain the probability ${\cal P}^{GeneTraj}[\vec{S}(0\leq t\leq T)]$ of this trajectory in the backward-generative dynamics. The hope is that the slight difference between $P_{*}(\vec{S}(T))$ and $P_{T}(\vec{S}(T))$ for large $T$ will produce probabilities ${\cal P}^{GeneTraj}[\vec{S}(0\leq t\leq T)]$ of generated-backward trajectories that are close to the probabilities ${\cal P}^{Traj}[\vec{S}(0\leq t\leq T)]$ of the forward trajectories, and will lead in particular to generated images $\vec{S}(0)$ at time $t=0$ with a probability distribution sufficiently close to the initial distribution $P^{ini}(\vec{S}(0))$.
The generated distribution $B^{Gene}_{t}(\vec{S}(t))$ at time $t$ is obtained from the application of the backward propagator ${\cal B}_{t,T}[\vec{S}(t)|\vec{S}(T)]$ of Eq. 7 to $P_{*}(\vec{S}(T))$
$B^{Gene}_{t}(\vec{S}(t))$ $\equiv\sum_{\vec{S}(T)}{\cal B}_{t,T}[\vec{S}(t)|\vec{S}(T)]P_{*}(\vec{S}(T))=P_{t}(\vec{S}(t))\sum_{\vec{S}(T)}P_{*}(\vec{S}(T))\frac{\langle\vec{S}(T)|{\cal W}^{(T-t)}|\vec{S}(t)\rangle}{P_{T}(\vec{S}(T))}$ (10)
$=\langle\vec{S}(t)|{\cal W}^{t}|P^{ini}\rangle\sum_{\vec{S}(T)}P_{*}(\vec{S}(T))\frac{\langle\vec{S}(T)|{\cal W}^{(T-t)}|\vec{S}(t)\rangle}{\langle\vec{S}(T)|{\cal W}^{T}|P^{ini}\rangle}$
and can be compared with the forward solution $P_{t}(\vec{S}(t))=\langle\vec{S}(t)|{\cal W}^{t}|P^{ini}\rangle$ that appears explicitly as prefactor before the sum over the configuration $\vec{S}(T)$.
For the special case $t=0$, the generated distribution of Eq. 10
$B^{Gene}_{0}(\vec{S}(0))$ $=\sum_{\vec{S}(T)}{\cal B}_{0,T}[\vec{S}(0)|\vec{S}(T)]P_{*}(\vec{S}(T))=P^{ini}(\vec{S}(0))\sum_{\vec{S}(T)}P_{*}(\vec{S}(T))\frac{\langle\vec{S}(T)|{\cal W}^{T}|\vec{S}(0)\rangle}{P_{T}(\vec{S}(T))}$ (11)
@@
can be compared with the initial distribution $P^{ini}(\vec{S}(0))$ that appears explicitly as prefactor before the sum over the configuration $\vec{S}(T)$.
II.1.5 Discussion
In conclusion, all the properties of the non-trivial backward dynamics can be rewritten in terms of the initial condition $P^{ini}$ and in terms of the forward propagator, in particular:
(i) at the level of a single time-step with the reweighting formula of Eq. 6 for the time-dependent generator $\langle\vec{S}(t)|{\cal W}^{B}_{t+1/2}|\vec{S}(t+1)\rangle$;
(ii) at the level of any finite-time-interval $[t_{1},t_{2}]\subset[0,T]$ with the backward propagator ${\cal B}_{t_{1},t_{2}}[\vec{S}(t_{1})|\vec{S}(t_{2})]$ of Eq. 7;
(iii) at the level of the probabilities of the whole trajectories $\vec{S}(0\leq t\leq T)$ on the time-window $[0,T]$ with the reweighting formula of Eq. 9.
For any time $t\in[0,T]$, the discrepancy between the generated distribution $B^{Gene}_{t}(\vec{S}(t))$ and the forward solution $P_{t}(\vec{S}(t))=\langle\vec{S}(t)|{\cal W}^{t}|P^{ini}\rangle$ can be analyzed via Eq. 10 that simplifies into Eq. 11 for the special case $t=0$ where one can compare the probability $B^{Gene}_{0}(\vec{S}(0))$ of the generated images with the initial distribution $P^{ini}(\vec{S}(0))$.
After this summary of the general framework, let us now focus on the explicit choice of the forward dynamics.